OpenStudy (anonymous):

Derivative of the following: F(x)= ((3x^2 -2)^2)/x^2

OpenStudy (amorfide):

okay to make this easier you want to bring the x² to the top when you do this, the power changes to negative so you get (3x²-2)²x^-2 multiply out of the brackets then differentiate as normal can you do this?

OpenStudy (cwrw238):

or you could use the product rule

OpenStudy (anonymous):

I'm still confused

OpenStudy (cwrw238):

are you familiar with chain rule and product rule ?

OpenStudy (anonymous):

Yes I am.

OpenStudy (anonymous):

Just not exactly when they are together

OpenStudy (cwrw238):

so applying these f'(x) = 6x*2(3x^2 - 2) * x^-2 + (3x^2 - 2)^2 * -2x^-3

OpenStudy (cwrw238):

product rule d(uv)/dx = u*dv/dx + v*(du/dx)

OpenStudy (cwrw238):

second thoughts - its probably a little easier to use the quotient rule!!

OpenStudy (cwrw238):

d(u/v)/ dx = [ v du/dx - u dv/dx] v^2 where u and v are functions of x f'(x) = x^2* 12x(3x^2 - 2) - (3x^2-2)^2 * 2x ---------------------------------- x^4 which needs to be simplified

OpenStudy (anonymous):

Hmmmmm not following

OpenStudy (cwrw238):

is it the differentiation of (3x^2 - 2)^2 that you are not following?

OpenStudy (cwrw238):

- you use chain rule to do this

OpenStudy (cwrw238):

do you use the f'(x) notation or dy/dx?

OpenStudy (anonymous):

F'(x)

OpenStudy (cwrw238):

ok chain rule states that (fg(x)' = f'(g)x * g'(x) here g(x) = 3x^2 - 2 and fg(x) = 3(x^2 - 2)^2 g'(x) = 6x right?

OpenStudy (anonymous):

Yes

OpenStudy (cwrw238):

and f'(g)x = 2(3x^2 - 2) so multiplying give 12x(3x^2 - 2)